An experiment with three outcomes has been repeated 50 times, and it was learned that E1 occurred 20 times, E2 occurred 12 times, and E3 occurred 18 times. Assign probabilities to the following outcomes for Ei , Ez and E3 . Round your answer to two decimal places. P(E1) P(E2) P(E3) What method did you use? Classical Relative Frequency Subjective Select
Answer #1

Answer

EVENT Frequency
E1 20
E2 12
E3 18

Total frequency = 20+12+18 = 50

We know that the probability is defined as the ratio of favorable outcome and total outcome

So, P(E1) = (favorable)/(total outcome) = 20/50 = 0.40

Similarly

P(E2) = (favorable)/(total outcome) = 12/50 = 0.24

And

P(E3) = (favorable)/(total outcome) = 18/50 = 0.36

This method is known as relative frequency method, i.e. in this method we use relative frequency to calculate the respective probability.